OpenStudy (anonymous):

what... |-2x + 3| = 15

OpenStudy (anonymous):

we know either \(-2x+3=15\) or \(-2x+3=-15\)

OpenStudy (anonymous):

solve both equations

OpenStudy (jhannybean):

and absolute value around the function means that the function has equidistant values from the left and the right In solving this you would have -2x + 3 = - 15 or -2x + 3 = 15 subtract 3 from both sides -2x +3 - 3 = - 15 -3 or -2x + 3 - 3 = 15 - 3 -2x = -18 or -2x = 12 divide by -2 -2x/-2 = -18/-2 or -2x/-2 = 12/-2 x = ? or x = ?